Suppose the simplified consolidated balance sheet shown below is for the entire chartered banking system. The banks' reserve ratio is 10 percent and the public does not wish to hold any cash balances.

Assets (1) Reserves 15,0000Securities 60,0000Loans 25,0000Liabilities (1) Bank Deposits 100,0000


a) What is the money multiplier?

Multiplier = 0


b) How much excess reserves does the chartered banking system have?

Excess Reserves = $0


c) What is the maximum additional amount the banking system might lend?

Maximum Amount = $0


d) Show in column 1 above how the consolidated balance sheet would look after the entire amount from part c) has been lent
Answer #1

Given,

Required reserve ratio = 10% = 0.10

a) Money multiplier = 1/Required reserve ratio = 1/0.10 = 10

b) Required reserves = Required reserve ratio * Deposits = 0.10*$1,000,000 = $100,000

Actual Reserves = $150,000

Excess Reserves = Actual Reserves - Required Reserves = $150,000 - $100,000 = $50,000

c) Maximum additional desposits that can be created due to the additional loans given = Excess reserves * Money multiplier = $50,000 * 10 = $500,000

Total Deposits = $1,000,000 + $500,000 = $1,500,000

Required reserves = 0.10 * $1,500,000 = $150,000

Loans = Total Deposits - Required Reserves - Securities = $1,500,000 - $150,000 - $600,000 = $750,000

Maximum additional loans the banking system might lend = $750,000 - $250,000 = $500,000

d)

Assets Liabilities
Reserves 150,000 Bank Deposits 1,500,000
Securities 600,000
Loans 750,000
Total 1,500,000 Total 1,500,000